OpenStudy (anonymous):

a manufacturer samples 100 wires for quality testing. four of he wires are found to be defective. if 750 wires are produces in one hour, how many should the manufacturer expect to be defective?

OpenStudy (anonymous):

@Hero

hero (hero):

Hint, if 4/100 wires are defective, then 1 out of every 25 wires are expected to be defective since: \[\frac{4}{100} = \frac{1}{25}\] Therefore solve: \[\frac{1}{25} = \frac{x}{750}\] To find the number of wires expected to be defective after one hour.

OpenStudy (anonymous):

sorry but my test is on maintenance i didnt finish it

OpenStudy (anonymous):

ill try my best tommorow

hero (hero):

Did you find the value of x?

OpenStudy (anonymous):

no its been a long day its late here now

OpenStudy (anonymous):

ill get the value of x tommorow

hero (hero):

Okay, good luck.

OpenStudy (anonymous):

thanks